OpenStudy (anonymous):

Giving medals!!!!!! Simplify.square root of 10 times square root of 8 square root of 18 4 square root of 5 square root of 80 None of the above

OpenStudy (jdoe0001):

hint: \(\bf \sqrt{x}\cdot \sqrt{y}\implies \sqrt{x\cdot y}\)

OpenStudy (jdoe0001):

same root, the radicands multiply each other

OpenStudy (anonymous):

so I multiply 10 and 8 ?

OpenStudy (jdoe0001):

well.... keeping the radical,yes what did you get?

OpenStudy (anonymous):

square root of 80

OpenStudy (jdoe0001):

ok so \(\large { \sqrt{10}\cdot \sqrt{8}\implies \sqrt{80} \\ \quad \\ {\color{brown}{ 80\to 2\cdot 2\cdot 2\cdot 2\cdot 5\to 2^2\cdot 2^2\cdot 5\to (2^2)^2\cdot 5}} \\ \quad \\ \sqrt{80}\implies \sqrt[{\color{blue}{ 2}}]{(2^2)^{\color{blue}{ 2}}\cdot 5}\implies ? }\)

OpenStudy (anonymous):

I'm not sure but, I thought it would be neither of them because I thought your not supposed to add the square roots.

OpenStudy (jdoe0001):

hmm we didn't add any of the roots

OpenStudy (anonymous):

multiply sorry

OpenStudy (jdoe0001):

you can, IF the root is the same in this case both are root 2, or square root, so they're both the same, thus they're multiplyable if the root were different, say \(\Large \sqrt[3]{10}\cdot \sqrt[5]{8}\) then we couldn't

OpenStudy (jdoe0001):

so \(\large { \sqrt{10}\cdot \sqrt{8}\implies \sqrt{80} \\ \quad \\ {\color{brown}{ 80\to 2\cdot 2\cdot 2\cdot 2\cdot 5\to 2^2\cdot 2^2\cdot 5\to (2^2)^2\cdot 5}} \\ \quad \\ \sqrt{80}\implies \sqrt[{\color{blue}{ 2}}]{(2^2)^{\color{blue}{ 2}}\cdot 5}\implies ? }\) what do you think?

OpenStudy (anonymous):

I'm not entirely sure about how to do this, so it's not the square root of 80 right?

OpenStudy (jdoe0001):

well... yes it's notice the 1st line but notice the red lines is 2 * 2 * 2 * 2 * 5 \(\ne\ 80?\)